# Words as Trigger Points in Social Media Discussions: A Large-Scale Case Study about UK Politics on Reddit

### v3

@@
Abstract
Political debates on social media sometimes flare up. From that moment on, users engage much more with one another; their communication is also more emotional and polarised. While it has been difficult to grasp such moments with computational methods, we suggest that trigger points are a useful concept to understand and ultimately model such behaviour. Established in qualitative focus group interviews to understand political polarisation (Mau, Lux, and Westheuser 2023), trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned. In the original studies, individuals show strong and negative emotional responses when certain triggering words or topics are mentioned. Our paper finds that these trigger points also exist in online debates. We examine online deliberations on Reddit between 2020 and 2022 and collect $>$100 million comments from subreddits related to a set of words identified as trigger points in UK politics. Analysing the comments, we find that trigger words increase user engagement and animosity, i.e., more negativity, hate speech, and controversial comments. Introducing trigger points to computational studies of online communication, our findings are relevant to researchers interested in affective computing, online deliberation, and how citizens debate politics and society in light of affective polarisation.
1 Introduction
@@
Online debates may suddenly escalate, and users intensely engage in emotional and confrontational deliberations. While these turning points are difficult to model and predict, they are essential moments in online communication. They mark the moment when a discussion is likely to get off the rails, and social media users stop debating with one another and instead debate against one another. Our study introduces trigger points as a useful concept to understand why and when social media users engage in such behaviour.
Trigger points, introduced by Mau, Lux, and Westheuser (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions. They open the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to adverse reactions, hate speech, and abusive language in online debates.
@@
Figure 1: Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
@@
Trigger points are related to how individuals perceive themselves in society. People have subjective expectations about acceptable behaviour, such as explicit knowledge of law-like rules and, more importantly, an implicit understanding of appropriate actions. These deeply rooted beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world. People feel triggered if these normative understandings are questioned. More formally, Mau, Lux, and Westheuser (2023, 246) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract. Trigger points jeopardise what individuals consider the fabric of society and their own role in it.” The authors suggest that trigger points can affect individuals through four mechanisms.
@@
  • Inequality. When expectations about equality might not be met. Either similar people are treated unequally, or, to the contrary, people who should be treated differently are treated the same. For example, social benefits might be at eye level with minimal wage earners.
@@
  • Slippery slope. There could be a fear of society developing norms about appropriate behaviour that goes in subjectively undesired directions and the impression that the individual might not have the power to do something against it. Claims such as “Opening the borders so that a country could be flooded with immigrants” is an example.
@@
  • Behavioural demands. When society may be perceived as demanding unreasonable behaviour. The role of pronouns is a point in case where the generic masculine pronouns “he/him” might trigger some, and the neutral pronoun “they/them” might trigger others.
Going against deeply held beliefs about individuals and their societal role causes discomfort. People respond to such triggers in an affective behavioural mode, increasing a debate’s intensity and emotionality.
In the social media context, we expect that this has two direct consequences (Figure 1). Trigger points lead to:
  • Hypothesis 1: more user engagement, which leads to more messages or comments.
  • Hypothesis 2: higher levels of animosity, which causes polarisation, negativity, anger, and hate.
@@
Our study seeks to establish the concept of trigger points in online communication. The research design and case selection closely follow this logic and examine whether and how individual words can act as trigger points in online debates. To this end, we conduct a large-scale case study and select five words that are ex-ante likely to work as trigger points in the context of UK politics. Specifically, we look into the words Rwanda, Brexit, NHS, Vaccine, Feminism. Collecting $>$100 million related Reddit comments, we use text analytics and NLP methods to understand whether these words lead to potentially negative, emotional, and harmful responses. The analysis builds upon a systematic comparison between a treatment and control group in three different scenarios:
  1. 1.
    Space: specific subreddits where the selected words trigger users.
  2. 2.
    Contextual: words that appear in a similar context to the target words but without the triggering component.
  3. 3.
    Time: periods identified as relevant for each specific trigger word.
We find that the trigger words cause higher engagement and more animosity, measured as an increase in a thread’s controversiality, negative sentiment, anger, and hate speech.
2 Related Work
Existing work studies how particular political words or phrases lead to more engagement (Brady et al. 2017; Hameleers et al. 2021) and posts with more emotion and animosity (Friedman et al. 2004; Martin et al. 2021; Schroth, Bain-Chekal, and Caldwell 2005; Rho and Mazmanian 2020) in various online settings. However, they lack a joint conceptual and theoretical framework—something we aim to provide.
Trigger points are fundamental to understanding polarisation and division in online discussions. Given that they are at the root of different conflictive behaviours later on, they help understand the causes of disagreements in online conversations (De Kock and Vlachos 2021) and their escalation (De Kock and Vlachos 2022; van der Meer et al. 2023) triggering explicit and implicit ad hominem attacks, hate speech, abusive or offensive language on social media (Talat and Hovy 2016; Davidson et al. 2017; Basile et al. 2019; Ousidhoum et al. 2019; Sap et al. 2019; ElSherief et al. 2021; Ocampo et al. 2023).
@@ 2 Related Work
Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al. 2019), Blodgett et al. (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model. Studying trigger points helps track the origins of bias, assess whether a social media post can lead to toxic comments (Dahiya et al. 2021), and how language affects polarisation on social media (Karjus and Cuskley 2024). It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon, Brady, and Van Bavel 2022).
Related to trigger points are dog-whistles—coded expressions that carry two meanings: a seemingly harmless one, and another one with a hidden hateful or provocative connotation that only insiders can decipher (Breitholtz and Cooper 2021). In social sciences, the concept of dog whistle has its roots in the study of racial politics in the United States, leading politicians to develop more subtle ways of appealing to racial prejudices (López 2013) and has evolved to encompass a broader range of phenomena, ranging from perpetuated systemic racism (Filimon and and 2024), to more broadly communication strategies of political ambiguity in identity politics(Tolvanen 2024), such as religion (Perry 2023) or immigration (Filimon 2016). Trigger points are similar in their equivocality and potential to evoke emotional responses. Conversely, using a dog whistle is conscious, but there is typically no intent when relying on trigger points. Moreover, while dog whistles are rhetorical devices to hide the real intention from some and create a rally-around-the-flag effect, trigger points are not hidden communication and incite a strong sense of otherness towards those covered by the trigger.
Research on the automation of identifying affective triggers in online discussions remains limited. For instance, Almerekhi et al. (2022) developed a framework to identify triggers for toxic behaviour within Reddit discussions. Similar work typically uses a restricted set of lexical features, primarily sentiment analysis, and often examines the level of engagement some topics may generate (Aldous, An, and Jansen 2023). Others, such as Beel et al. (2022), explicitly investigate highly controversial topics like abortion, climate change, and gun control, attempting to predict contentiousness within discussions, often heavily relying on Reddit comment upvote ratios as a proxy for contentiousness. Finally, some studies focus on online conversational tensions by predicting conversational derailment, considering the overall progression of a conversation rather than specific terms as causes (Zhang et al. 2018; Hessel and Lee 2019).
3 Methodology
To empirically demonstrate that individual words can trigger online debates, we look into most likely cases in the context of UK politics. Based on a curated set of likely candidate words for the political context in the UK (see Section 3.1), we test whether individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity. Our goal is to report behavioural mechanisms similar to those observed by Mau, Lux, and Westheuser (2023) when conducting interviews to analyse the perceived polarisation. If we could not find these effects in these intuitive cases, we would falsify the theory for online contexts.
3.1 Term selection
We selected National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism. These terms are associated with high-profile public debates, policy changes, and social movements in the UK. Hence, they typically provoke significant public engagement and emotional responses. We expect them to cause the intended reactions in British social media users. Table 1 indicates mechanisms through which the terms may trigger individuals.
@@ 3.1 Term selection
Table 1: Trigger words and their respective trigger mechanisms (described in the introduction).
Rwanda.
@@
Until late 2021, Rwanda was just the name of any other African country. However, in early 2022, the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership. [^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible With this agreement, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist 2023).
Brexit.
@@
The UK’s departure from the European Union has been controversial since its inception. The division among the population was already evident during the tightly contested 2016 referendum(Arnorsson and Zoega 2018). Additionally, misinformation and disinformation have intensified these debates (Höller 2021; Bruno, Lambiotte, and Saracco 2022). Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from Eastern Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
NHS.
@@
Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the healthcare system (The Guardian 2022; The Independent 2024). Public opinion is divided, with some supporting that the NHS receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick, Charlesworth, and Taylor 2015). This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
Vaccine.
@@
Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years. Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell 2021). Moreover, the proliferation of misinformation and conspiracy theories has further fuelled public division (Hayawi et al. 2022; Garett and Young 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
Feminism.
@@
Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz 2019). Some social media influencers have built a large audience by sharing and promoting harmful content against women. This has led to the formation of online echo chambers (Farrell et al. 2019) where such views are amplified and propagated. Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope, and unreasonable behavioural demands.
@@
Table 2: Subreddits selected for each trigger word.
@@
Table 3: Percentage of comments per thread before (Be) and after (Af) each trigger appearance (Hypothesis 1). Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Contextual, comparisons. Bold scores indicate statistically significant results. Detailed distribution of comments can be seen in Appendix B, see Figure 8.
3.2 Data collection
@@
Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al. 2020). We searched for exact matches of specific keywords in comments from 2019 to 2022 and then collected the entire discussion thread for each extracted comment. [^2] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume. We excluded comments in languages other than English, identified as such using a fastText-based language identifier (Bojanowski et al. 2017). Comments marked as “[deleted]” by the API were also removed. Following this, only threads that contained comments before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point.
As a final filtering step, we limited our analysis to comments 30 minutes before and 30 minutes after the first mention of the triggering word. Previous research has shown that Reddit comments typically receive the majority of interactions within a short window after the comment is created (Thukral et al. 2018). Moreover, early responses are often important in shaping the direction of the conversation (Singer et al. 2016; Horne, Adali, and Sikdar 2017). Considering the fast-paced nature of Reddit, the 30-minute window filters out unrelated responses and ensures the focus of our analysis remains on the most relevant data [^3] Table 7 in Appendix B shows the total number of comments, threads, and comments collected for each trigger word and the period they were gathered..
3.3 Research design
@@
We identify the causal effect of trigger points with the difference-in-difference estimator (DiD), which systematically compares a treatment with a control group (Angrist and Pischke 2009; Card and Krueger 1994). It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group. In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms and those in the control group to be much less triggered by the respective words. There should be a more substantive behavioural change for the treatment group than for the control group.
Control groups
@@
We systematically compare treatment and control groups along three different dimensions. [^4] Identifying the causal effect in three completely unrelated comparisons helps address concerns about eventual endogeneity.
1. Space
@@
Discussions on Reddit are structured in different fora, also called subreddits. Aiming to investigate the effects a trigger word might have on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 2). We compare them with discussions in a control group, i.e., the rest of Reddit with many non-domain specific subreddits.
2. Contextual
@@
We also compare trigger words with terms that appear in a similar context, but that does not go against the more profound underlying transgressions responsible for trigger responses in similar contexts. To members of the subreddits in Table 2, using an explicit trigger word or its more harmless control term should make a real difference. We define related control group words and contrast woman to Feminism/Feminist, Red Cross to NHS, EU/European Union to Brexit, surgery to vaccine, and names of Rwanda’s neighbouring countries [^5] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia. to Rwanda. Treatment and control groups are from the same subreddits in the same year.
3. Temporal
@@
In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time. For “Rwanda”, we compare discussions in subreddits likely to spark reactions in 2020 to discussions in 2022. For Vaccine we compare threads from 2019 (before the COVID-19 pandemic) to 2022’s (post COVID-19 pandemic). Again, we expect that users from the subreddits in Table 2 are triggered by those only in 2022 but not in the respective earlier years. Treatment and control groups share the same subreddits but are from different years.
Statistical analysis

### v4

# Causal Effects of Trigger Words in Social Media Discussions: A Large-Scale Case Study about UK Politics on Reddit

@@
Abstract.
Political debates on social media often escalate quickly, leading to increased engagement as well as more emotional and polarised exchanges. Trigger points (Mau et al., 2023) represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is being questioned. Analysing Reddit discussions, we examine how trigger points shape online debates and assess their impact on engagement and affect. Our analysis is based on over 100 million comments from subreddits centred on a predefined set of terms identified as trigger words in UK politics. We find that mentions of these terms are associated with higher engagement and increased animosity, including more controversial, negative, angry, and hateful responses. These results position trigger words as a useful concept for modelling and analysing online polarisation.
Affective polarization, Causal effect, Computational social science, Hate speech, Reddit, Trigger words, UK politics
* Equal contribution
journalyear: 2026copyright: ccconference: 18th ACM Web Science Conference; May 26–29, 2026; Braunschweig, Germanybooktitle: 18th ACM Web Science Conference (WebSci ’26), May 26–29, 2026, Braunschweig, Germanydoi: 10.1145/3795766.3799750isbn: 979-8-4007-2504-3/2026/05ccs: Applied computing Sociologyccs: Information systems Social networksccs: Computing methodologies Discourse, dialogue and pragmatics
1. Introduction
@@
Online debates may suddenly escalate, prompting users to engage in emotional and confrontational deliberations. While this exact moment is challenging to model and predict, it is essential in online communication. It signals when discussions derail and users shift from debating to attacking one another. Our study introduces trigger words as a useful concept for understanding why and when social media users engage in such behaviour—which is central to accounting for hate speech and other abusive online interactions.
@@
Figure 1. Consequences of trigger words. Hypothesis 1: Trigger words cause more engagement; Hypothesis 2: Trigger words cause more animosity.
@@
Trigger words are related to how individuals perceive themselves in society. People hold internal models of what counts as acceptable behaviour. These models include explicit, rule-like knowledge, e.g., formal norms or laws. More importantly, they also contain implicit, hard-to-articulate expectations about which actions are appropriate in a given context. Such deeply rooted beliefs about what is fair and what is unfair constitute a moral compass fundamental to navigating the social world. People feel triggered if these normative understandings are questioned (Figure 1). More formally, Mau et al. (2023) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract. Trigger points jeopardise what individuals consider the fabric of society and their own role in it.” In our study, we investigate to what extent specific words in online discussions can behave as such trigger points. Trigger points can affect individuals through four mechanisms (Mau et al., 2023):
@@
Inequality. When expectations about equality might not be met. Either similar people are treated unequally, or, to the contrary, people who should be treated differently are treated the same. For example, social benefits might be at eye level for minimum-wage earners.
@@
Slippery slope. People may fear a society developing norms about appropriate behaviour that go in subjectively undesired directions. They may feel they lack the power to do anything about it. Claims such as “Opening the borders so that a country could be flooded with immigrants” are an example.
@@
Behavioural demands. A society may be perceived as demanding unreasonable behaviour. The role of pronouns is a case in point where the generic masculine pronouns “he/him” might trigger some, and the neutral pronoun “they/them” might trigger others.
Because these trigger points cause emotional discomfort, people respond in an affective-behavioural mode, increasing the debate’s intensity and emotionality. In the social media context, this has two direct consequences (Figure 1). Trigger words lead to: (i) Hypothesis 1: more user engagement, which leads to more messages or comments; and (ii) Hypothesis 2: higher levels of animosity, which causes polarisation, negativity, anger, and hate.
@@
The goal of our study is to highlight and establish the concept of trigger words in scholarly work on online communication. Our research design and case selection closely follow this logic: We want to show that even the mention of a single trigger word can escalate online debates. To this end, we select five words that are ex-ante likely to work as trigger words in the context of UK politics (Rwanda, Brexit, NHS, Vaccine, Feminism). Collecting >>100 million in related Reddit comments, we utilise text analytics and classifiers to determine whether these words elicit potentially negative, emotional, or harmful responses. The analysis builds on a systematic comparison between a treatment and control group in three different scenarios: (1) Space: specific subreddits where the selected words are likely to trigger users; (2) Contextual: words that appear in a similar context to the trigger words but without the triggering component; and (3) Temporal: periods identified as relevant for each specific trigger word.
2. Related Work
@@
Existing work studies how particular political words or phrases lead to more engagement (Brady et al., 2017; Hameleers et al., 2021) and posts with more emotion and animosity (Friedman et al., 2004; Martin et al., 2021; Schroth et al., 2005; Rho and Mazmanian, 2020) in various online settings. However, they lack a joint conceptual and theoretical framework—something we aim to provide. Trigger words are fundamental to understanding polarisation and division in online discussions. Given that they are at the heart of different conflictive behaviours later on, they help understand the causes of disagreements in online conversations (De Kock and Vlachos, 2021) and their escalation (De Kock and Vlachos, 2022; van der Meer et al., 2023) triggering explicit and implicit ad hominem attacks, hate speech, abusive or offensive language on social media (Talat and Hovy, 2016; Davidson et al., 2017; Basile et al., 2019; Ousidhoum et al., 2019).Similarly, given a broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al., 2019), Blodgett et al. (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model. Studying trigger words helps track the origins of such behaviour, assess whether a social media post can lead to toxic comments (Dahiya et al., 2021), and how language affects polarisation on social media (Karjus and Cuskley, 2024). It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon et al., 2022).
Trigger words resemble dog whistles—coded expressions with dual meanings, one benign and one provocative to insiders (Breitholtz and Cooper, 2021). Originating in US racial politics (López, 2013), the dog whistle concept covers systemic racism (Filimon and and, 2024), political ambiguity in identity discourse (Tolvanen, 2024), religion (Perry, 2023), and immigration (Filimon, 2016). Dog whistles are used deliberately to obscure intent. Trigger words, however, are not covert but evoke strong affective reactions and heighten perceived otherness.
@@
Table 1. Our trigger words and their respective trigger mechanisms (see introduction).
Research on automating the identification of affective triggers in online discussions remains limited. For instance, Almerekhi et al. (2022) developed a framework to identify triggers for toxic behaviour within Reddit discussions. Similar work typically uses a restricted set of lexical features, primarily sentiment analysis, and often examines the level of engagement some topics may generate (Aldous et al., 2023). Others, such as Beel et al. (2022), explicitly investigate highly controversial topics, including abortion, climate change, and gun control, attempting to predict contentiousness within discussions. They often heavily rely on Reddit comment upvote ratios as a proxy for contentiousness. Finally, some studies focus on online conversational tensions by predicting conversational derailment, considering the overall progression of a conversation rather than specific terms as causes (Zhang et al., 2018; Hessel and Lee, 2019).
3. Methodology
To empirically demonstrate that individual words can act as triggers in online debates, we look into the most likely cases in the context of UK politics. Our goal is to report behavioural mechanisms similar to those observed by Mau et al. (2023) offline, i.e., (1) more engagement and (2) animosity. If we could not find these effects in these intuitive cases, we would falsify the theory for online contexts.
3.1. Term selection
We selected National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism. These terms are associated with high-profile public debates, policy changes, and social movements in the UK. Hence, they typically provoke significant public engagement and emotional responses, i.e., the intended reactions in British social media users. Table 1 suggests mechanisms through which the terms may trigger individuals.
@@ 3.1. Term selection
Rwanda. Until late 2021, Rwanda was simply one of many African countries in the eyes of prevailing UK public discourse. However, in early 2022, the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership. [^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible With this agreement, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist, 2023).
@@
Brexit. The UK’s departure from the European Union has been a contentious issue since its inception. The division among the population was already evident during the tightly contested 2016 referendum (Arnorsson and Zoega, 2018). Additionally, misinformation and disinformation have intensified these debates (Höller, 2021; Bruno et al., 2022). Considerations about treating EU migrants differently from the national population (inequality), norm transgressions of “criminal” migrants, or opening the gate to “unlimited” immigration make Brexit a potential trigger word.
@@
NHS. Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the healthcare system (The Guardian, 2022; The Independent, 2024). Some support the claim that the NHS receives adequate funding, while others argue that it wastes resources or that its services are in decline (Gershlick et al., 2015). This trigger word is related to going against considerations about unequal treatment.
@@
Vaccine. Recently, public vaccination programs, particularly those against COVID-19, have caused significant debate. Post-roll-out, debates have shifted to focus on the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell, 2021). Moreover, the proliferation of misinformation and conspiracy theories has fueled public division (Hayawi et al., 2022; Garett and Young, 2021), particularly in social media.
@@
Feminism. The anonymity provided by social media often emboldens users to express misogyny (Barker and Jurasz, 2019). Some influencers have built a large audience by sharing and promoting harmful content against women. This has led to online echo chambers (Farrell et al., 2019) where such views are amplified and propagated. We include Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope, and unreasonable behavioural demands.
@@
Table 2. Subreddits for each trigger word.
@@
Table 3. Percentage of comments per thread before (Be) and after (Af) each trigger appearance (Hypothesis 1). Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space, and Contextual comparisons. Bold scores indicate statistically significant results.
3.2. Data collection
@@
We collected English-language Reddit comments via the Pushshift API (Baumgartner et al., 2020). Searching for exact matches of specific keywords in comments from 2019 to 2022, we collected the entire discussion thread for each extracted comment. [^2] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume. We excluded comments in languages other than English using a fastText-based language identifier (Bojanowski et al., 2017) and also removed comments that the API marked as “[deleted]”. Only threads with comments both before and after the target keywords were retained to enable analysis of pre- and post-trigger differences.
As a final step, we included only comments posted within 30 minutes before and after the first mention of the trigger word. This window captures the most relevant responses, given that Reddit interactions peak quickly and early replies often shape the conversation’s direction (Thukral et al., 2018; Singer et al., 2016; Horne et al., 2017).
3.3. Research design
@@
We estimate the causal effect of trigger words using the difference-in-differences (DiD) estimator, which systematically compares a treatment group with a control group (Angrist and Pischke, 2009; Card and Krueger, 1994). It calculates the causal effect of an intervention as the difference in the treatment group before and after the intervention minus the difference in the control group before and after the intervention. In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms, whereas those in the control group are expected to be much less likely to be triggered by the exact words.
3.3.1. Control groups
@@
Discussions on Reddit are organised in different fora (subreddits). We systematically compare the treatment and control groups across three dimensions. [^3] Identifying the causal effect in three completely unrelated comparisons helps address concerns about eventual endogeneity.
@@
1. Space: To investigate the effects of trigger words on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 2). We compare them with discussions in a control group, i.e., the rest of Reddit, which includes many non-domain-specific subreddits.
@@
2. Contextual: We also compare trigger words to similar terms that lack the deeper transgressive meaning. For users in the subreddits listed in Table 2, the difference between a trigger word and a milder control term should make a difference. We define control terms by pairing, for example, woman with feminism/feminist, Red Cross with NHS, EU/European Union with Brexit, surgery with vaccine, and Rwanda with its neighbouring countries [^4] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia.. Treatment and control data come from the same subreddits and year.
@@
3. Temporal: In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time. For “Rwanda”, we compare discussions in subreddits likely to spark reactions in 2020 with those in 2022. For Vaccine, we compare threads from 2019 (pre-COVID-19 pandemic) to 2022 (post-COVID-19 pandemic). Again, we expect that users from the subreddits listed in Table 2 were triggered only in 2022, not in earlier years. The treatment and control groups share the same subreddits but differ in year.
3.3.2. Statistical analysis